# How to Use Large Language Models for Text Coding: The Case of Fatherhood Roles in Public Policy Documents

### v2

@@ Abstract
Recent advances in large language models (LLMs) like GPT-3 and GPT-4 promise automation with better results and less programming, opening up new opportunities for text analysis in political science. In this study, we evaluate LLMs on three original coding tasks of non-English political science texts, and we provide a detailed description of a general workflow for using LLMs for text coding in political science research. Our use case offers a practical guide for researchers looking to incorporate LLMs into their research on text analysis. We find that, when provided with detailed label definitions and coding examples, an LLM can be as good as or even better than a human annotator while being much faster (up to hundreds of times), considerably cheaper (costing up to 60% less than human coding), and much easier to scale to large amounts of text. Overall, LLMs present a viable option for most text coding projects. Our software allows to easily use LLMs as annotators and it is publicly available at https://github.com/lorelupo/pappa.
@@ 1 Introduction
Political science is among the social sciences that have most eagerly embraced automatic text analysis methods grimmer2021machine (22). Indeed, in many downstream research applications, such as data collection, topic classification, and especially content analysis, political scientists have found natural language processing (NLP) techniques to hold significant potential in terms of cost-saving and scaling capacity. Various applications (text scores, topic models, etc.) have been introduced and, in many cases, adapted to the specific needs of the field slapin2008scaling (43, 21, 38). Yet, despite their promise, data scarcity (the lack of sufficient high-quality training data and the cost and time associated with creating them) and a lack of technical expertise have thus far prevented NLP methods from being widely implemented—both in political science and in the social sciences more generally.
@@
Recently, a new generation of large language models (LLMs) has changed the cost-efficiency calculus for political scientists interested in the possibilities of NLP techniques. LLMs, including the notorious GPT series and Bard, can not only produce coherent texts almost indistinguishable from those of human writers, they can also (at least partially) automate many tasks—including coding text for analysis. LLMs need only a few examples written in plain language to learn a coding scheme, a property called “few-shot learning” brown2020language (8). This is possible because the model has a general concept of language and only needs to be specialized on the task at hand, a significant improvement over both traditional statistical text classification approaches and the most modern transfer learning ones laurer_van2023_trans (31), which require thousands of hand-coded data points. In addition, web interfaces like ChatGPT and APIs allow the use of LLMs with little programming knowledge. [^1] These models are called “closed” because their inner workings are not known. Several “open” language models exist; however, they typically require some amount of programming knowledge to use. In this paper, we focus on GPT models with a web interface. In the future, companies like HuggingFace will likely provide access to many open models via an API. These properties make LLMs attractive for political scientists who work with text but might not have the computational knowledge or resources required for other approaches. LLMs represent a shift in the potential scale, cost-efficiency, and broad accessibility of text analysis. It is likely to change the type and volume of research done, as well as the methods and the volume of data used (and hence the reliability of the results). This is the case not only for the technical fields in which LLMs were developed (computer science, AI, NLP, and machine learning) but also in political science, higher education, and various other disciplines.
@@
In this paper, we use a concrete political science project to illustrate the workflow of text annotation with LLMs. Specifically, we code “fatherhood roles” in Swedish political discourse (see Section 3 for details). The different fatherhood roles we define are theoretical constructs with subtle differences from one another, making the coding task harder than many typical NLP coding tasks like sentiment analysis, where human annotators typically have high agreement because of the net difference between the labels. We define three different, interdependent coding tasks of varying difficulty, which we then use to code an original Swedish policy-related document dataset from 1993–2021. We annotate the data with trained human coders, GPT-3, and GPT-4.222We do not cover open-source models like LLaMa touvron2023llama (44), as they require specialized programming knowledge to use. However, our open-source software includes the possibility to use them as annotators. We also experiment with and guide the reader through some relevant choices in the LLM-coding methodology (Section 4.5): joint versus disjoint coding of multiple interdependent tasks, level of detail in the prompt description of labels, zero-shot versus few-shot learning, and the order of coding examples provided in few-shot learning. For each of these settings, we compare the performance of the language models to that of human coders along the various tasks, evaluating the quality, cost, and speed of both human coders and machines. Importantly, our evaluation is not subject to “data contamination” golchin2023time (18) because we can guarantee that the LLMs have not encountered the tasks we are evaluating during their training. Contamination arises when models are trained on popular, open-source text datasets that may include the very tasks on which they are evaluated, and potentially even the specific text examples that are present in their test set. Such exposure during training leads to models essentially memorizing correct responses, resulting in inflated performances on the “contaminated” test sets. This phenomenon affects popular datasets across various scientific fields, including political science, casting doubt on the reported performance of models’ evaluated on them. While current LLMs might have read the Swedish documents collected in our dataset during their training phase, the coding tasks are surely novel to them and cannot have been memorized.
To summarize, our main contributions are:
  1. 1.
    We find that individual coders have widely differing label outcomes, which do not always agree with others.
  2. 2.
    Evaluating two versions of GPT (3 and 4) on three political science coding tasks, using different prompting strategies, we find the best way to use the models is to provide them with comprehensive label definitions and several coding examples (i.e., few-shot learning).
  3. 3.
    In this setting, GPT-3 does not quite reach human performance, but it is comparable. GPT-4, in contrast, consistently outperforms human coders, agreeing more with each individual than the humans do with each other. Both LLMs are orders of magnitude faster than any human coder.
  4. 4.
    We find that jointly performing three coding tasks on the same texts results in comparable performance to performing them separately; the former approach is also approximately three times cheaper and faster.
2 Language Models in Political Science Research
AI, with its promise of harnessing an abundance of data, has been eagerly adopted within political science grimmer2021machine (22). Using various NLP and machine learning (ML) techniques, “computer assistants” have been used to read and accurately assess political texts in a fraction of the time it would take human research assistants to perform the same tasks. These techniques, which heavily rely on ML models benoit_2020text (7), have allowed for sophisticated analyses, from descriptive tasks such as understanding the delegation of power in the EU anastasopoulos_2020understanding (4) to more interpretative assignments, such as topic classification wang_2023_trans (46) and the surveillance of vast amounts of text to detect and flag political conflict with considerable accuracy hu_2022conflibert (27). Computer assistants have proven their significant potential for political science.
However, considerable thresholds for outsourcing downstream research tasks to computer assistants remain. Typically, ML models require large amounts of manually annotated data to train for the assignment. And once trained, they are not easily adapted to another assignment. Combined with the technical expertise required to maneuver these models, this data scarcity creates considerable barriers to implementing computer assistants, begging the question of whether switching from human to machine is worth the effort.
The data scarcity issue has triggered research dedicated to transfer learning osnabrugge_2023trans (34, 31): training the model on a dataset similar to what you want to label, thus relaxing the requirements for original annotated data.
The recent launch of Generative Pretrained Transformers (GPTs) can lower the barrier to entry even further. Developed by OpenAI, GPT language models radford2018improving (36, 37, 8, 32) can be accessed through web interfaces such as ChatGPT, OpenAI’s Playground, or programmatically through an API. In 2020, OpenAI presented GPT-3 brown2020language (8), which boasts 175 billion parameters and underwent training on a vast dataset of over 45 terabytes of text. This setup makes it one of the biggest generative language models ever developed, only surpassed by a handful of other LLMs—including its younger sibling GPT-4, the most recent and powerful model deployed by OpenAI openai2023gpt4 (32). GPTs have several practical advantages over other LLMs, particularly over a family of language models commonly used in coding tasks: BERT models.
The first is that GPT can be queried through an intuitive interface and does not require any fine-tuning (the neural network remains unchanged for whatever coding task), only proper instructions (prompting). Most BERT-style models must be loaded on your PC or server. The necessary software needs to be installed and then run on thousands of examples to fine-tune the model to the coding task at hand, a task that typically requires powerful computers with GPUs.
The second advantage is that, unlike all but a very few traditional language models, GPT can work with multiple languages. Moreover, GPT can handle a long set of instructions and examples. In contrast, BERT models need the classification task to be expressed in one sentence or a little more, making them unsuitable for classification tasks requiring accurate description. Furthermore, BERT models cannot simultaneously code a text across multiple tasks, as we show to be possible with GPT.
In economic research, (korinek_language_2023, 29, ) demonstrates 25 scenarios where LLMs like GPTs can automate tasks like ideation, analysis, and software development. Empirical studies have explored LLMs’ ability to mimic human behaviors in surveys and experiments. Notably, (argyle_out_2023, 5, ) revealed GPT-3’s capacity to replicate diverse survey responses, and (aher_using_2023, 2, ) successfully used LLMs to simulate human subjects in research, reproducing key social science findings. Extending this, (horton_large_2023, 26, ) showed GPT-3’s effectiveness in economic gaming simulations, while dillion_can_2023 (13) found GPT-3.5’s moral assessments align with human judgments. Furthermore, rosenbusch2023accurate (40) discovered GPT-3’s accuracy in generating hypotheses about social phenomena, highlighting the potential of LLMs in research automation.
These pioneering investigations merely scratched the surface of LLMs’ potential to streamline research processes. Unsurprisingly, LLMs show great potential in automating text coding assignments ornstein2022train (33). In this article, we contribute to this line of research by illustrating how language models can effectively automate and apply a novel theoretical framework to code original empirical data.
@@ 3 The Example Case Study
As a case study for the proposed methodology, we investigate how fathers are discussed in Swedish policy documents, i.e., government investigations, propositions, motions, and other legislative texts. The background of the study is the emergence of a “new father”—the idea that societal discourse is becoming less dominated by a traditional fatherhood role, and more open to other alternatives—and the debate, in academia and elsewhere, about whether and when such a shift occurred, the content of the shift, and its consequences. The study uses data from Sweden over the period 1993–2021. The substantive research question—whether fathers have been assigned a more agentic fatherhood role over time in these documents—will be elaborated on in a forthcoming article.
Stereotypical fatherhood roles in texts are more or less always analyzed using qualitative interpretive analyses graf2023post (19, 20). This means that political science has developed valuable theoretical and normative concepts and understands specific texts and cases empirically dermott2015more (12, 23, 35). However, there is a dramatic under-usage of quantitative text analyses and few possibilities for concluding whether fathers are recognized differently in different types of documents, over different periods and contexts, and by different types of message senders. For example, it is highly unclear whether societal norms around fatherhood are changed via elected representatives (so that norm change is found in motions, speeches, or the like) or via public officials and administrations (public investigations), or decided by who ends up in government (propositions, government regulations, etc.). Furthermore, the absence of quantitative measures in the field renders it unclear whether the different fatherhood roles are common or uncommon in these texts. [^3] For a recent study using experimental designs to study fatherhood roles, see (scheifele2021representations, 42, ) Thus, this case study represents a political science research area where large-scale automatic text coding is needed.
@@ 3.1 Data
We used publicly available data from the Swedish parliament (Riksdagen). [^4] https://www.riksdagen.se/sv/dokument-och-lagar/riksdagens-oppna-data/ We collected all public documents from 1993 to 2021, converted them to raw text format, and lowercased them. We selected only sentences referencing fathers (i.e., containing the Swedish words pappa, pappor, fäder, fader, far). We added further checks to ensure the term actually referred to a father: for example, the word “far” can also mean “go” or “travel” if used as a verb. We used a part-of-speech tagger to filter out those instances. The final data set contains 1,910 sentences. Data collection and filtering details are reported in B.
@@ 4 Methodology
In this section, we present the main steps researchers should follow in an LLM coding task after collecting their data:
  1. 1.
    Coding paradigm: deciding what the labels are.
  2. 2.
    Validation set: hand-code a few hundred examples to evaluate the LLM’s performance.
  3. 3.
    Language model: choose one, or compare multiple models before selecting the most suitable.
  4. 4.
    Prompt: construct the textual input to the model, including a concrete instruction, a description of the labels, and a few hand-coded examples (which should not be part of the validation set).
  5. 5.
    Prompt engineering: fine-tune the prompt by trying out different options. For each option, let the LLM code the validation set and evaluate its label predictions against human labels.
  6. 6.
    Coding the entire corpus by querying the LLM with the best-performing prompt obtained from the previous step.
@@ 4.1 Coding paradigm
After an initial qualitative analysis of our dataset, we classified each of the Swedish texts on fatherhood along three tasks, each represented by a varying number of suitable labels:
  1. 1.
    Type of paternal involvement in the family. Possible labels for this task are “passive,” “active negative,” “active positive caring,” “active positive daring,” “active positive other,” and “not applicable.”
  2. 2.
    Explicitness of the description, denoting the linguistic means by which the father’s role is expressed in the sentence. Possible labels: “implicit” and “explicit.”
  3. 3.
    Normativeness of the description, denoting whether the sentence expresses a factum or a desired outcome. Possible labels: “descriptive” and “ideal.”
The three labeling tasks target different aspects of the texts (content vs. form), are interconnected, and are hierarchical. If the first task is “not applicable”, the other two tasks are irrelevant and thus not coded. B presents a detailed description of each label.
@@ 4.2 Validation set
Three authors independently coded the same random subset of 350 sentences, applying the coding paradigm described above. This validation set allows us to establish a baseline of human performance. It gives us a direct comparison set for the coding choices by the LLMs. This step is optional but highly recommended, as it is needed to measure the performance of the LLMs and to carry out the prompt engineering step to maximize performance.
4.3 Language Model
For most of our experiments, we used the GPT-3 model from OpenAI via their API. [^5] https://platform.openai.com/docs/api-reference/introduction. Specifically, we used the text-davinci-003 language model to classify texts into our predetermined labels. [^6] https://platform.openai.com/docs/models Due to the model’s language pretraining, it is sophisticated enough to learn the labeling task from just a prompt, the components of which will be discussed in the following section. Following the best prompting strategy identified in the prompt engineering step, we also evaluated the most powerful (and most expensive) OpenAI model available to date: GPT-4. However, even for that model, the overall cost is moderate (see Section 5.3).
@@ 4.4 Prompt
For GPT, we used the exact same instructions and label definitions as would be used in the codebook for human coders. However, how they are provided to the model differs along some dimensions.
Instructions
Language models need clear instructions that explicitly describe the goal of the task. Instructions can be formulated in any of the languages understood by the model. However, performance might vary according to the languages used. We used English instructions, even though the text coded by the model is in Swedish. In fact, in preliminary experiments we found that Swedish instructions alongside Swedish texts resulted in inferior performances compared to English instructions. The most likely reason for this behavior is that GPT-3 is mainly trained on English data brown2020language (8), and, therefore, English is the language it understands best. This finding suggests that GPT should be a stronger coder of English texts than non-English ones, as corroborated in other recent studies ahuja2023mega (3, 14).
The instruction we used for coding all tasks reads as follows:
@@
Labels
We next provided the codebook to the language model prompt. Besides providing the model with the list of labels defined in the previous coding paradigm step (Section 4.4), the model can also benefit from a description of these labels. Depending on the labeling task, the labels might be already self-explanatory (such as for task 2) or unclear (such as for task 1). As we will see in the experimental section, in the latter case an exhaustive description, including ample formulation of the details and hints to identify relevant texts, usually works better than a synthetic description, which in turn is better than no description at all. Our codebook, including labels and their definitions, can be found in B.1. In the experiments, we simply copy-pasted this codebook into the prompt after the instruction.
Example sentences
Following the codebook, we provided GPT-3 with 15 example sentences from our corpus that we manually labeled with the three desired labels. GPT-3 used these examples to learn how sentences map to labels and apply this mapping to the texts given to it. It is important to include examples for all of the labels, and that these examples are not part of the validation set. We randomly sampled these examples from our corpus, excluding the validation set. The examples used in our experiments are reported in C.1.
The procedure of providing a handful of examples in the prompt is called few-shot learning. The selection of these examples can significantly impact the coding performance of the language model. For this reason, we recommend measuring the model’s performance on the validation set when provided with different sets of examples. Note that each example is counted against our word budget, with more words meaning higher costs (see Section 5.3).
Alternatively, we can perform zero-shot learning by leaving out all examples and providing only the instructions and label descriptions. This is much faster, less prone to incorporate human biases, and potentially cheaper, but known to be potentially less accurate.
@@ 4.5 Prompt engineering
To optimize GPT’s coding performance, we evaluated results from different prompting combinations, varying the following features of the codebook and the example sentences:
@@
Indicates whether the model is prompted to code a single task at a time, or the three tasks simultaneously, for each text that is provided in input. In the former case, the model outputs a single label for each text, while in the latter it generates three, one for each task. For example, in Table 1, we prompted the model to code task 1 alone, except in the last two rows, were we ask to code all tasks. The idea behind the second option is to use shared information in the hierarchical coding scheme, and to make the coding process more efficient, as the model does not have to be run three times on the same instance to provide a label for each task.
@@
Indicates different degrees of explanation of the labels provided to the model, from a long and comprehensive explanation to a concise one to none. We experimented with all three variants to code task 1 (type of paternal involvement), where labels have complex semantics. In contrast, we coded the second (explicitness) and third (normativeness) tasks with self-explanatory labels. Further explanation is in these cases redundant and hence not provided to the model.
@@
5 Evaluation
We evaluate the automatic coding along three variables: quality, speed, and cost.
5.1 Quality
One advantage of letting GPT-3 label text instead of humans is the higher chance of getting consistent coding over time. Humans are known to be inconsistent (they become hungry, get a text message, need to go to the bathroom) and often disagree with their previous assessments when queried again abercrombie-etal-2023-temporal (1). They also frequently and systematically disagree with other human coders, even if they have received the same training and instructions. Coding using GPT-3 should be more consistent, as long as temperature is kept to 0 (this parameter controls how “creative” GPT’s answers are).
To investigate the reliability of humans compared to GPT, we first compare the performance of three human coders (all authors of the paper) to each other on the validation set. We evaluate their coding along three measures: chance-corrected agreement (Cohen’s $\kappa$), raw agreement (Raw), and predictive performance (F1). Cohen’s $kappa$, a measure of inter-annotator agreement, is usually between 40 and 60 landis1977measurement (30). Less-than-perfect results for the human coders do not necessarily signal bad training or bad coders but usually indicate how difficult a task is and the absence of a ground truth. To measure the raw agreement of an annotator, say “Human 1,” with the others, we use the other human annotators’ set of labels as the gold standard and measured how often Human 1’s label predictions matched them on average. The predictive performance is calculated similarly, but accounting for class imbalances (i.e., gold label “passive” being more present in the test set than the label “active_negative”). The results of this evaluation can be found in the first block of rows of Table 1, 2, and 3, where the first, second, and third labeling tasks are evaluated, respectively.
Besides providing the performance of each human annotator, we also measure their average performance (“Humans AVG”).
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 49.30 61.29 57.22
Human 2 n.a. n.a. n.a. 51.02 63.86 57.60
Human 3 n.a. n.a. n.a. 51.65 64.29 57.54
Humans AVG n.a. n.a. n.a. 50.66 63.15 57.45
GPT-3 1 none 0 13.41 27.71 22.73
GPT-3 1 short 0 20.89 36.38 34.27
GPT-3 1 none 15 36.53 51.05 44.22
GPT-3 1 short 15 44.50 58.48 49.40
GPT-3 1 long 15 49.36 62.57 54.92
GPT-3 1,2,3 long 15 48.34 62.48 54.26
GPT-4 1,2,3 long 15 54.05 66.38 61.70
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 26.74 54.86 50.13
Human 2 n.a. n.a. n.a. 26.55 58.29 48.61
Human 3 n.a. n.a. n.a. 29.75 59.14 52.33
Humans AVG n.a. n.a. n.a. 27.68 57.43 50.36
GPT-3 2 none 0 14.66 42.76 34.72
GPT-3 2 none 15 22.91 53.14 48.54
GPT-3 1,2,3 none 15 25.91 59.14 50.07
GPT-4 1,2,3 none 15 31.97 63.62 53.49
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 41.64 66.57 61.83
Human 2 n.a. n.a. n.a. 37.00 67.14 57.94
Human 3 n.a. n.a. n.a. 44.17 71.14 63.20
Humans AVG n.a. n.a. n.a. 40.94 68.28 60.99
GPT-3 3 none 0 7.35 40.29 29.94
GPT-3 3 short 0 7.96 36.86 32.00
GPT-3 3 none 15 39.75 68.76 60.97
GPT-3 3 short 15 44.08 70.86 64.25
GPT-3 1,2,3 short 15 41.85 70.48 62.67
GPT-4 1,2,3 short 15 40.55 71.81 61.43
The task is sufficiently difficult even for trained, motivated, and expert human coders. The average Cohen’s $\kappa$ indicates “substantial” agreement, though on the lower end of that designation. The takeaway is that annotating fatherhood roles is tricky.
Using zero-shot learning (i.e., zero examples) results in bad performance across the board. Longer, comprehensive label definitions help whenever they are relevant, i.e., for tasks one and three. However, they cannot match the improvement we see with 15-shot learning, even without an explanation of the labels. The combination of few-shot learning and long, comprehensive label definitions results in the best performance, comparable to that of human coders and sometimes even outperforming them, as was the case for our third task.

### v3

# Towards Human-Level Text Coding with LLMs: The Case of Fatherhood Roles in Public Policy Documents

@@ Abstract
Recent advances in large language models (LLMs) like GPT-3.5 and GPT-4 promise automation with better results and less programming, opening up new opportunities for text analysis in political science. In this study, we evaluate LLMs on three original coding tasks involving typical complexities encountered in political science settings: a non-English language, legal and political jargon, and complex labels based on abstract constructs. Along the paper, we propose a practical workflow to optimize the choice of the model and the prompt. We find that the best prompting strategy consists of providing the LLMs with a detailed codebook, as the one provided to human coders. In this setting, an LLM can be as good as or possibly better than a human annotator while being much faster, considerably cheaper, and much easier to scale to large amounts of text. We also provide a comparison of GPT and popular open-source LLMs, discussing the trade-offs in the model’s choice. Our software allows LLMs to be easily used as annotators and is publicly available at https://github.com/lorelupo/pappa.
@@ 1 Introduction
Political science is among the social sciences that have most eagerly embraced automatic text analysis methods (grimmer2021machine). Indeed, in many downstream research applications, such as data collection, topic classification, and especially content analysis, political scientists have found natural language processing (NLP) techniques to hold significant potential in terms of cost-saving and scaling capacity (slapin2008scaling; grimmer2010bayesian; roberts2016model). Yet, despite their promise, data scarcity (the lack of sufficient high-quality training data and the cost and time associated with creating them) and a lack of technical expertise have thus far prevented NLP methods from being widely implemented—both in political science and the social sciences more generally.
@@
Recently, a new generation of large language models (LLMs) has changed the cost-efficiency calculus for political scientists interested in the possibilities of NLP techniques. LLMs, including the renowned GPT series, can not only produce coherent texts almost indistinguishable from those of human writers, but they can also automate many tasks—including coding text for analysis. This is possible because the model has a general concept of language and only needs to be instructed on how to solve the task at hand, a significant improvement over both traditional statistical text classification approaches and the most modern transfer learning ones (laurer_van2023_trans), which still require thousands of hand-coded data points. In addition, APIs like OpenAI’s or the Hugging Face Hub allow the use of LLMs with little programming knowledge. These properties make LLMs attractive for political scientists who work with text but might not have the resources to manually annotate thousands of examples and the computational knowledge to run supervised approaches.
@@
In this paper, we use a concrete political science project to study text annotation with LLMs. Specifically, we define three different, interdependent coding tasks of varying difficulty about “fatherhood roles” in Swedish political discourse (see Section 2 for details). While several research projects have already shown the LLM’s prowess in text annotation, few - if any - have investigated their performance in a typical political science setting involving a minority language (like Swedish) with legal or political jargon and complex labels (see ollion2023chatgpt for an overview). Indeed, the different fatherhood roles we define are theoretical constructs with subtle differences from one another, making the coding task harder than many typical NLP coding tasks like sentiment analysis, where human annotators typically have a high intercoder agreement. Moreover, the novelty of our annotation task offers a robust test of the LLMs’ generalizability, as the possibility that the LLM has memorized the task during its training (data contamination) can be ruled out. Finally, evaluating LLMs on complex tasks where a ”ground truth” does not exist requires benchmarking against multiple human annotators. One element stands out in our contribution, compared to the existing literature: we acknowledge this lack of consensus by assessing the LLMs against not one, but three trained political scientists. In our work, we evaluate the performance of two renowned GPT models (GPT-3 and GPT-4) and three Open-Source (OS) LLMs, experimenting with some relevant choices in the LLM-coding methodology (Section 3.5): joint versus disjoint coding of multiple interdependent tasks, level of detail in the prompt description of labels, zero-shot versus few-shot learning, and the order of coding examples provided in few-shot learning. For each of these settings, we compare the performance of the language models to that of three human coders along the various tasks.
Our main findings are: 1) LLMs annotation is optimal when they are treated as human coders: providing them with an exhaustive codebook of label definitions and several coding examples; 2) Following this prompting strategy, some LLMs perform on par or outperform human coders, even in complex annotation tasks involving a minority language, linguistic specificity, and complex labels; 3) The popular GPT-4 model is at the frontier of annotation performance in the most difficult task, while OS LLMs catch up on the easier tasks; 4) Jointly performing three coding tasks on the same texts results in comparable performance to performing them separately (approximately three times cheaper and faster).
@@ 2 The Example Case Study
As a case study for the proposed methodology, we investigate the occurrence of different predefined father roles in Swedish policy documents. The task is derived from a real-world research question on changing societal norms in Sweden from a forthcoming paper and incurs a large degree of interpretation on the annotators as it relies on complex abstract constructs, that were defined specifically for this research question. In this sense, the task is theoretically challenging, and the low agreement between human annotators (see 4) indicates that annotators disagree on the ”right” label in more than just a few cases. So far, the literature has evaluated LLMs on coding tasks that were either theoretically simpler (such as the political polarity of a sentence), or based on English texts, where current LLMs are typically the strongest, or both (see ollion2023chatgpt for an overview). Our task is representative of many cases encountered in political science research, involving complex labels and minority languages. Moreover, the novelty of out coding task guarantees that our evaluation is not subject to “data contamination” (golchin2023time) since it is improbable that the LLMs have encountered the tasks we are evaluating during their training. Contamination arises when models are trained on popular, open-source text datasets that may include the very tasks on which they are evaluated, and potentially even the specific text examples that are present in their test set. Such exposure during training leads to models essentially memorizing correct responses, resulting in inflated performances on the “contaminated” test sets. This phenomenon affects popular datasets across various scientific fields, including political science, casting doubt on the reported performance of models evaluated on them. While current LLMs might have read the Swedish documents collected in our dataset during their training phase, the coding tasks are surely novel to them and cannot have been memorized.
@@ 2.1 Data
We used publicly available data from the Swedish parliament (Riksdagen) [^1] https://www.riksdagen.se/sv/dokument-och-lagar/riksdagens-oppna-data/ including government investigations, propositions, motions, and other legislative texts. We collected all public documents from 1993 to 2021, converted them to raw text format, and lowercased them. We selected only sentences referencing fathers (i.e., containing the Swedish words pappa, pappor, fäder, fader, far). We added further checks to ensure the term actually referred to a father: for example, the word “far” can also mean “go” or “travel” if used as a verb. We used a part-of-speech tagger to filter out those instances. The final data set contains 1,910 sentences. Data collection and filtering details are reported in B.
@@ 3.1 Coding paradigm
After an initial qualitative analysis of the dataset, the theoretically derived labels were applied to a small (16) sample of the data. We classified each of the Swedish texts on fatherhood along three tasks, each represented by a varying number of suitable labels. The three labeling tasks target different aspects of the texts (content vs. form), are interconnected, and hierarchical. If the first task is “not applicable”, the other two tasks are irrelevant and thus not coded. B presents a detailed description of each label.
Task 1 - Type of paternal involvement
in the family. Possible labels for this task are “passive,” “active negative,” “active positive caring,” “active positive daring,” “active positive other,” and “not applicable.”
Task 2 - Explicitness of the description
, denoting the linguistic means by which the father’s role is expressed in the sentence. Possible labels: “implicit” and “explicit.”
Task 3 - Normativeness of the description
, denoting whether the sentence expresses a factum or a desired outcome. Possible labels: “descriptive” and “ideal.”
@@ 3.2 Validation set
A validation set was created (350 sentences) to establish a baseline of human performance and provides a direct comparison set for the LLMs’ coding choices. Measuring LLMs performance allows one to optimize the prompt provided to them and to choose the most suitable LLM. To build a validation set, researchers should provide human annotators with a codebook describing the task and the labels, and providing some labelled examples. Three authors independently coded the same random subset of 350 sentences following our codebook, which can be found in B.1.
@@ 3.3 Prompt
Similarly to humans, LLMs need to be instructed to annotate text. In our experiments we evaluated different prompting strategies, showing that the best way to prompt LLMs is by providing them with the same codebook that was built for humans. Here we describe its main components: the instruction, the description of the labels, and some labelled examples. The codebook usually starts with an instruction defining the task. The instruction we used for all of our three coding tasks reads as follows: [^2] Note that we used English instructions, even though the text coded by the model is in Swedish. In fact, in preliminary experiments, we found that Swedish instructions alongside Swedish texts resulted in inferior performances compared to English instructions.
@@
The list of labels defined in the coding paradigm step (Section 3.1), should then follow. Each label should also be accompanied by an exhaustive description of its meaning. In our case, the labels of the first task needed longer descriptions, as the definitions of fatherhood roles are more abstract and complex than the labels of the other two tasks. Labelled examples should also be provided to the coders, especially highlighting cases of ambiguity. We randomly sampled example sentences from our corpus (excluding the validation set) and annotated them. We discussed and agreed on the labels assigned to the examples.
3.4 Language Models
Many options exist today when it comes to LLMs, both open source (OS) and proprietary, like the GPTs. Every LLM has distinct costs and performances on different tasks and different languages. For a new coding task, performances are hard to know in advance. Evaluating the most promising LLMs on the validation set allows to for an informed choice. We experimented with two of the best proprietary LLMs to date, i.e., the GPT-3 model text-da-vinci-003 and GPT-4 (openai2023gpt4) from OpenAI333via their API https://platform.openai.com/docs/api-reference/introduction, and two of the best OS models: Llama-2 (the -13b-chat-hf version; touvron2023llama) and Mixtral-8x7b (the -Instruct-v0.1 version. Alongside, we also tested a less popular OS model which was fine-tuned on Nordic languages: GPT-SW3 (the -20b-instruct version; ekgren2023gptsw3).
@@ 3.5 Prompt engineering
To optimize LLMs’ coding performance, different prompting combinations can to be evaluated on the validation set. In our case, we experimented with the following elements of the codebook:
@@
Indicates whether the model is prompted to code a single task at a time or the three tasks simultaneously for each text that is provided in input. In the former case, the model outputs a single label for each text, while in the latter it generates three, one for each task. For example, in Figure 1, we prompted the model to code all the three tasks together in the column ”multi-task”. The motivation is to allow the model to use shared information in the hierarchical coding scheme and to make the coding process more efficient, as the model does not have to be run three times on the same instance to provide a label for each task.
@@
Indicates different degrees of explanation of the labels provided to the model. From no description, to a short description, to a long and comprehensive one (as the labels’ description provided in the codebook). We experimented with all the three variants to code task 1 (type of paternal involvement), where labels have complex semantics. In contrast, we coded the second (explicitness) and third (normativeness) tasks with self-explanatory labels. Further explanation is in these cases redundant and hence not provided to the model.
@@
4 Empirical Results
To investigate humans inter-reliability, we compare the predictive performance (macro-F1) of each of the three human coders against the labels of the other two. We then plot the humans’ scores average (green dashed line) and standard deviation (green shade) in Figures 1, 2, and 3 for tasks 1, 2, and 3, respectively. In the same Figures we plot the performance average and standard deviation of each prompt configuration and model evaluated against humans’ labels.
Figure 1: Average coding agreement with (other) human coders on task 1: type of paternal involvement (6 labels).
Figure 2: Average coding agreement with (other) human coders on task 2: explicitness of the description (2 labels).
Figure 3: Average coding agreement with (other) human coders on task 3: normativeness of the description (2 labels).
The average F1 score for humans ranges between 60 and 50 for the tasks. Less-than-perfect results for the human coders do not necessarily signal bad training or bad coders but usually indicate how difficult a task is, and the absence of a ground truth. The takeaway is that annotating fatherhood roles is tricky.
4.0.1 Prompt Engineering
The first bars on the left of each figure represent the performance of GPT-3 with different prompting strategies. In the first bar, GPT-3 is prompted to annotate tests in a zero-shot learning (zero labelled examples) and without any description of the labels beside their list. This prompt results in bad performance across the board. Switching to few-shot learning (second bar in each Figure) significantly improves performances, even in the absence of any description of the labels. Including labels’ descriptions in relevant tasks (tasks 1 and 3; the labels for task 2 are self-explanatory) further improves results, especially when descriptions are exhaustive, i.e., long descriptions for the task 1, short descriptions for task 3. The combination of few-shot learning and comprehensive label definitions results in the best performance, comparable to that of human coders. In other words, GPT-3 performs best when prompted with the same codebook that is provided to human coders.